# Resonances and resonance expansions for $\delta$-interactions on the half-space

### v2

@@ 1. Introduction
Schrödinger operators $H$ on unbounded domains, besides a rich and much developed spectral theory, are characterized by the existence of resonances, i.e. poles of the meromorphic continuation in the lower complex half-plane of the resolvent $(H-z^{2})^{-1}$. Loosely speaking, resonances behave as a sort of complex eigenvalues; associated to them there are resonance functions, solving formally the eigenvalue equation but only locally square integrable. The use of resonances is common in Physics and their importance is due to the fact that when $H$ is the generator of a Schrödinger or wave dynamics, they correspond to states decaying in time (at least for certain time regimes), the slower ones (and so the more physically meaningful) being the nearest to the real axis. Among other reasons, resonances are of interest because the solution of the Wave or Schrödinger equation propagating on non-compact domains can be expressed, asymptotically for large times and for initial data localized in space, as an expansion based on resonance functions [1, 33, 22], as an alternative to the expansion based on the spectral theorem. For these and other reasons, it is desirable to have as precise information as possible about the location and properties of resonances of self-adjoint operators. While a lot of rigorous results have been accumulated along the years about general qualitative properties of resonances and their several applications in Physics, Dynamical Systems theory and Geometry (see the treatise [22] for an encyclopedic and recent account), precise and quantitative information also for specific classes of operators is however rare. In this paper we want to begin the analysis of the resonances of the class of self-adjoint singular perturbations of the Laplacian called $\delta$-interactions or point interactions and considered on domains of $\mathbb{R}^{n},\ n=2,3$. There is a huge literature on point interactions (see [2] for the bibliography until 2005, still growing). Their interest is due mostly to the fact that, in many case, they represent exactly solvable models in Quantum Mechanics and that they are suitable scaling limit of more realistic, non solvable models. However, they have been studied almost exclusively on the whole space $\mathbb{R}^{n},\ n=1,2,3$. Only a handful of papers treats the case of bounded domains (see [12, 26]) of hyperbolic space (see the recent [21], including also point interactions on spheres), while the case of unbounded domains not coinciding with $\mathbb{R}^{3}$ has been treated essentially only in the geometries of infinite tubes and layers, especially relevant to the subject of quantum waveguides (see the book [25]). As regards resonances, the laplacian with a single point interaction in $\mathbb{R}^{n},\ n=1,2,3$ has at most a single resonance, a strong difference with standard Schrödinger operators $H=-\Delta+V$. On the contrary, $N$-point delta interactions have infinitely many resonances. This has been recognized originally in [4], and later the distribution of complex resonances has been investigated again by Albeverio and Karabash [7, 6, 5] and by Lipovský and Lotoreichik [34]. The special case of real resonances in dimension 3 is discussed in [35]. Properties of the resolvent in dimension two are instead studied in [15], but the analysis of resonances of point interactions in dimension two is still rather poor. The one dimensional case is the best known, and see in particular the recent papers [18] and [39] and references therein. We stress that the treatment of the one dimensional case is greatly simplified by the fact that the $\delta$ distribution can be considered as a form bounded perturbation of the laplacian and can be treated, somewhat formally but safely, as a standard potential. This is by no means the case in higher dimension, where the point interaction cannot be properly interpreted as a kind of Schrödinger operator "$-\Delta+\delta$".
@@
As regards resonances, little is known when the domain considered is an unbounded subset of $\mathbb{R}^{n}$. We mention in one dimension the half-line with a delta potential and a Dirichlet boundary conditions (called Winter model in the physical literature) has received some attention ([17, 39] and references therein) and in the higher dimensional case some examples of waveguides (see [25]). Here we consider probably the case of the half-space with Dirichlet or Neumann boundary conditions on the boundary, and a point interaction at a generic position $y$. The plan of the paper is the following. In Section 2 point interactions on the half-space are introduced, giving their complete definition and relevant properties. Because a rather standard construction is involved in the rigorous definition, we omit proofs of Proposition 2.1 and Proposition 2.2. We stress that these model cases do not reduce to any combinations of point interactions in $\mathbb{R}^{3}$ (see Remark 2.5 and Remark 3.8). Then, in Section 3 we obtain the explicit calculation of all the resonances of the considered models, here with full details and relevant remarks. The explicit knowledge of the set of resonances allows, in Section 4, to give results about their counting in spherical neighborhoods of given radius in the complex plane. It should be noticed that the distribution does not resembles the analogous one for smooth potentials in dimension three. This different behavior was already noticed in [34] for several point interactions in $\mathbb{R}^{3}$ and it is here confirmed for a domain not coinciding with $\mathbb{R}^{3}$. In Section 5 we study the semiclassical asymptotic of resonances, which parallels the one dimensional case treated recently in [17]. Finally, in Section 6 the expansion in resonances is deduced, both for the wave and Schrödinger evolution. More precisely, we give the expansion in resonances of the flow for the wave equation and of the propagator for the Schrödinger operator of a point interaction in the half-space. The resonance expansion for point interactions was previously considered in models containing point interactions in [1] for arrays of point interactions in $\mathbb{R}^{3}$ and in [11] where a quantized model of matter and radiation introduced in [37, 36] is studied. As a possible alternative to the direct analysis given here we notice that resonances in domains with point interactions could fit well in the black box description developed by Sjostrand and Zworski ([42, 41]). Black box analysis of resonances and the extension of the results obtained in the present paper to domains more general than the half-space, as the analysis of the two dimensional case, seem to be interesting tasks and they will be the object of future research.
@@ 2. Preliminaries: Point Interaction on the half-space
The half-space is $\mathbb{R}_{+}^{3}=\Set{x=(x_{1},x_{2},x_{3})\in\mathbb{R}^{3}}{x_{3}>0}$ and its boundary is +$\partial\mathbb{R}^{3}_{+}=\left\{x=(x_{1},x_{2},x_{3})\in\mathbb{R}^{3}\ |\ x_{3}=0\right\}$.
We denote with $G_{z,y}^{0}$ the Green’s function centered at $y\in\mathbb{R}^{3}$ for the operator $-\Delta-z^{2}$ on the whole space
@@
The Green function centered at $y\in\mathbb{R}^{3}_{+}$ of the operator $-\Delta-z^{2}$ on the half-space $\mathbb{R}_{+}^{3}$ with Dirichlet or Neumann boundary conditions is given respectively by
@@
By a simple symmetry argument it follows the well-known fact that
@@
where $\overline{y}=(y_{1},y_{2},-y_{3})$ denotes the reflected of the point $y$ with respect to the plane $x_{3}=0$.
Moreover we will denote with $-\Delta^{\textup{D}}$ and $-\Delta^{\textup{N}}$ the free Laplacian on $\mathbb{R}_{+}^{3}$ with Dirichlet and Neumann boundary conditions respectively on +$\partial\mathbb{R}^{3}_{+}$, considered as self-adjoint operators on $L^{2}(\mathbb{R}^{3}_{+})$. These operators have the respective domains
$D(-\Delta^{\textup{D}})$ $=H^{2}(\mathbb{R}^{3}_{+})\cap H^{1}_{0}(\mathbb{R}^{3}_{+})$
$D(-\Delta^{\textup{N}})$ $=H^{2}(\mathbb{R}^{3}_{+})\ .$
The resolvent set of $-\Delta^{\textup{D}}$ and $-\Delta^{\textup{N}}$ is still $\mathbb{C}\setminus[0,+\infty)$.
Let us fix an arbitrary $y\in\mathbb{R}_{+}^{3}$, where the singular interaction is placed. A straightforward calculation allows to check that the symmetric non self-adjoint operators
$-\Delta^{\textup{D}}$ with $\ \ \text{with domain}\ \ \mathring{D}(-\Delta^{\textup{D}})=\left\{\psi\in L^{2}(\mathbb{R}^{3}_{+})\ \text{s.t.}\ \ \psi\in D(-\Delta^{\textup{D}}),\ \psi(y)=0\right\}$
$-\Delta^{\textup{N}}$ with $\ \ \text{with domain}\ \ \mathring{D}(-\Delta^{\textup{N}})=\left\{\psi\in L^{2}(\mathbb{R}^{3}_{+})\ \text{s.t.}\ \ \psi\in D(-\Delta^{\textup{N}}),\ \psi(y)=0\right\}$
have both defect indices $(1,1)$. So, by the Von Neumann-Kreĭn theory, they admit a one parameter family of self-adjoint extensions. Their definition and properties are given in the following propositions. Their proof is a rather standard adaptation of the analogous proof for the case of $\mathbb{R}^{3}$ (see for example the classical text [2]) and therefore it is omitted.
@@
Let $-\Delta_{\alpha,y}^{\textup{D}}$ be the one parameter family of self-adjoint extensions of the symmetric non self-adjoint operator $-\Delta^{\textup{D}}$ with domain $\mathring{D}(-\Delta^{\textup{D}})=\left\{\psi\in L^{2}(\mathbb{R}^{3}_{+})\ \text{s.t.}\ \ \psi\in D(-\Delta^{\textup{D}}),\ \psi(y)=0\right\}$.
Then, the following representation holds for the operator domain, operator action and resolvent of $-\Delta_{\alpha,y}^{\textup{D}}$ (where it is understood that $\operatorname{Im}z>0$):
$D\Bigl{(}-\Delta_{\alpha,y}^{\textup{D}}\Bigr{)}$ =$=\Set{\psi\in L^{2}\Bigl{(}\mathbb{R}_{+}^{3}\Bigr{)}}{\,\exists q\in\mathbb{R}\,\textup{s.t }\psi=\varphi^{z}+qG_{z,y}^{\textup{D}},\,\varphi^{z}\in D(-\Delta^{\textup{D}}),\,\varphi^{z}(y)=\Gamma_{\alpha,y}^{\textup{D}}q}$
@@
(2.1) $(-\Delta_{\alpha,y}^{\textup{D}}-z^{2})^{-1}\psi$ $=(-\Delta^{\textup{D}}-z^{2})^{-1}\psi+(\Gamma_{\alpha,y}^{\textup{D}})^{-1}(-\Delta^{\textup{D}}-z^{2})^{-1}\psi(y)G_{z,y}^{\textup{D}},\quad\forall\psi\in L^{2}(\mathbb{R}_{+}^{3})$
@@
In particular, $\alpha=+\infty$ corresponds to $-\Delta^{\textup{D}}$.
@@
Let $-\Delta_{\alpha,y}^{\textup{N}}$ be the one parameter family of self-adjoint extensions of the symmetric non self-adjoint operator $-\Delta^{\textup{N}}$ with domain $\mathring{D}(-\Delta^{\textup{N}})=\left\{\psi\in L^{2}(\mathbb{R}^{3}_{+})\ \text{s.t.}\ \ \psi\in D(-\Delta^{\textup{N}}),\ \psi(y)=0\right\}$. Then, the following representation holds for the operator domain, operator action and resolvent of $-\Delta_{\alpha,y}^{\textup{N}}$ (where it is understood that $\operatorname{Im}z>0$):
$D\Bigl{(}-\Delta_{\alpha,y}^{\textup{N}}\Bigr{)}$ =$=\Set{\psi\in L^{2}\Bigl{(}\mathbb{R}_{+}^{3}\Bigr{)}}{\,\exists q\in\mathbb{R}\,\textup{s.t }\psi=\varphi^{z}+qG_{z,y}^{\textup{N}},\,\varphi^{z}\in D(-\Delta^{\textup{N}}),\,\varphi^{z}(y)=\Gamma_{\alpha,y}^{\textup{N}}q}$
@@
(2.2) $(-\Delta_{\alpha,y}^{\textup{N}}-z^{2})^{-1}\psi$ $=(-\Delta^{\textup{N}}-z^{2})^{-1}\psi+(\Gamma_{\alpha,y}^{\textup{N}})^{-1}(-\Delta^{\textup{N}}-z^{2})^{-1}\psi(y)G_{z,y}^{\textup{N}},\quad\forall\psi\in L^{2}(\mathbb{R}_{+}^{3})$
@@
In particular, $\alpha=+\infty$ corresponds to $-\Delta^{\textup{N}}$.
@@
The generic domain element of a point interaction is decomposed in the sum of a regular part ($\varphi^{z}$) belonging to the domain of the laplacian with relevant boundary conditions and a singular part ($qG_{z,y}^{\textup{D}}$ or $qG_{z,y}^{\textup{N}}$) proportional to the Green’s function for the given domain and boundary condition. The decomposition is complemented by a relation linking the regular part at $y$ and the coefficient $q$ of the singular part, playing the role of a further boundary condition at the singular point $y$. The latter contains the parameter $\alpha$ which specifies the particular self-adjoint extension of the free laplacian among the whole family of its self-adjoint extensions. This pattern is typical of point interactions in their various incarnations. We add that the physical interpretation of $\alpha$ is related to the scattering length of the interaction.

### v3

@@ 1. Introduction
Schrödinger operators $H$ on unbounded domains, besides a rich and much developed spectral theory, are characterized by the existence of resonances, or more precisely continuation resonances, by which we mean the poles of the meromorphic continuation in the lower complex half-plane (in suitable topologies) of the resolvent $(H-z^{2})^{-1}$ (see the seminal paper [52] for the rigorous introduction of the concept). Loosely speaking, resonances behave as a sort of complex eigenvalues; associated to them there are resonance functions, solving formally the eigenvalue equation but only locally square integrable. Among the many reasons, resonances are of interest because the solution of the wave or Schrödinger equation propagating on non-compact domains can be expressed, asymptotically for large times and for initial data localized in space, as an expansion based on resonance functions [6, 39, 28], as an alternative to the expansion based on the spectral theorem. For these and other reasons, it is desirable to have as precise information as possible about the location and properties of resonances of self-adjoint operators. While a lot of rigorous results have been accumulated along the years about general qualitative properties of resonances and their several applications in Physics, Dynamical Systems and Geometry (see the treatise [28] for an encyclopedic and recent account), precise and quantitative information also for specific classes of operators is however rare. In this paper we want to begin the analysis of the continuation resonances of the class of self-adjoint singular perturbations of the Laplacian called $\delta$-interactions or point interactions on subdomains of $\mathbb{R}^{n},\ n=2,3$. There is a huge literature on point interactions (see [3] for the bibliography until 2005, still growing). Their interest is due mostly to the fact that, in many cases, they represent exactly solvable models in Quantum Mechanics and that they are suitable scaling limit of more realistic, non solvable models. However, they have been studied almost exclusively on the whole space $\mathbb{R}^{n},\ n=1,2,3$. Only a handful of papers treats the case of bounded domains (see [16, 32]) of hyperbolic space (see the recent [27], including also point interactions on spheres), while the case of unbounded domains not coinciding with $\mathbb{R}^{3}$ has been treated essentially only in the geometries of infinite tubes and layers, especially relevant to the subject of quantum waveguides (see the book [31]). As regards resonances, as it is well known, the Laplacian with a single point interaction in $\mathbb{R}^{n},\ n=1,2,3$ has at most a single resonance, a strong difference with typical Schrödinger operators $H=-\Delta+V$ (see for example [19]). On the contrary, $N$-point delta interactions have infinitely many resonances. This has been recognized originally in [4], and later the distribution of complex resonances has been investigated again by Albeverio and Karabash [7, 8, 9] and by Lipovský and Lotoreichik [40]. The special case of real resonances in dimension 3 is discussed in [42]. Properties of the resolvent in dimension two are instead studied in [22], but the analysis of resonances of point interactions in dimension two is still rather poor. The one dimensional case is the best known, see in particular the recent papers [24] and [47] and references therein. We stress that the treatment of the one dimensional case is greatly simplified by the fact that the $\delta$ distribution can be considered as a form bounded perturbation of the Laplacian and can be treated, somewhat formally but safely, as a standard potential. We mention in one dimension the half-line with a delta potential and a Dirichlet boundary conditions, which is called the Winter model in the physical literature and has received some attention in [23, 47] (see also the references therein). This is by no means the case in higher dimensions, where the point interaction cannot be properly interpreted as a kind of Schrödinger operator $\Delta+\delta$". As regards resonances, little is known when the domain considered is an unbounded subset of $\mathbb{R}^{n}$. In the higher dimensional case we refer to some examples concerning waveguides (see [31]).
@@
Here we consider the case of the half-space with Dirichlet or Neumann boundary conditions on the boundary, and a point interaction at a generic position $y$. The plan of the paper is the following. In Section 2 point interactions on the half-space are introduced, giving their complete definition and relevant properties. Because a rather standard construction is involved in the rigorous definition, we postpone the proofs of Proposition 2.1 and Proposition 2.2 to the Appendix. We stress that these model cases do not reduce to any combinations of point interactions in $\mathbb{R}^{3}$ (see Remark 2.5 and Remark 3.8). Then, in Section 3 we obtain the explicit computation of of all the resonances of the considered models, here with full details and relevant remarks. The explicit knowledge of the set of resonances allows us to give in Section 4 a result about the number of resonances in complex discs of radius $R$ as $R\to\infty$ . It should be noticed that the distribution does not resemble the analogous one for smooth radial potentials for the whole space. This different behavior was already noticed in [40] for several point interactions in $\mathbb{R}^{3}$ and it is here confirmed for a domain not coinciding with $\mathbb{R}^{3}$ (see remark 4.2 for more details on this). In Section 5 we study the semiclassical asymptotic of resonances, which parallels the one dimensional case treated recently in [23]. Finally, in Section 6 the expansion in resonances is deduced, both for the wave and Schrödinger evolution. More precisely, we give the expansion in resonances of the flow for the wave equation and of the propagator for the Schrödinger operator of a point interaction in the half-space. The resonance expansion for point interactions was previously considered in models containing point interactions in [6], where an array of point interactions in $\mathbb{R}^{3}$ is studied, and in [14], where a quantized model of matter and radiation introduced in [44, 43] is studied. Finally, we notice that a possible alternative to the direct analysis of continuation resonances given here is by means of the black box scattering theory developed by Sjostrand and Zworski in [49, 50]. Black box analysis of resonances and the extension of the results obtained in the present paper to domains more general than the half-space, as well as the analysis of the two dimensional case, seem to be interesting tasks and they will be the object of future research.
@@ 2. Preliminaries: Point Interaction on the half-space
[…]
We denote with $G_{z,y}^{0}$ the Green’s function centered at $y\in\mathbb{R}^{3}$ for the operator $\Delta-z^{2}$ on the whole space
@@
The Green function centered at $y\in\mathbb{R}^{3}_{+}$ of the operator $\Delta-z^{2}$ on the half-space $\mathbb{R}_{+}^{3}$ with Dirichlet or Neumann boundary conditions is given respectively by
@@
By a simple symmetry argument one obtains that
@@
where $\overline{y}=(y_{1},y_{2},-y_{3})$ denotes the reflection of the point $y$ with respect to the plane $x_{3}=0$.
Moreover we will denote with $\Delta^{\textup{D}}$ and $\Delta^{\textup{N}}$ the free Laplacian on $\mathbb{R}_{+}^{3}$ with Dirichlet and Neumann boundary conditions respectively on $\partial\mathbb{R}^{3}_{+}$, considered as self-adjoint operators on $L^{2}(\mathbb{R}^{3}_{+})$. These operators have the respective domains
$D(-\Delta^{\textup{D}})$ $=H^{2}\left(\mathbb{R}_{+}^{3}\right)\cap H^{1}_{0}\left(\mathbb{R}_{+}^{3}\right)$
$D(-\Delta^{\textup{N}})$ $=\left\{\psi\in H^{2}(\mathbb{R}_{+}^{3}):\left.\frac{\partial\psi}{\partial x_{3}}\right\rvert_{x_{3}=0}=0\right\}.$
where in the Neumann case the boundary value of the partial derivative has to be intended in the sense of trace (see for closely related framework [11] and for the specific case of the half-space [12]). Let us fix an arbitrary $y\in\mathbb{R}_{+}^{3}$, the place where the singular interaction is placed. We consider the symmetric non self-adjoint restrictions of the Dirichlet and Neumann Laplacians
$-\mathring{\Delta}_{y}^{\textup{D}}$ $\ \ \text{with domain}\ \ D(-\mathring{\Delta}_{y}^{\textup{D}})=\left\{\psi\in D(-\Delta^{\textup{D}}),\ \psi(y)=0\right\}$
$-\mathring{\Delta}_{y}^{\textup{N}}$ $\ \ \text{with domain}\ \ D(-\mathring{\Delta}_{y}^{\textup{N}})=\left\{\psi\in D(-\Delta^{\textup{N}}),\ \psi(y)=0\right\}.$
These restrictions are well defined because, by Sobolev embedding, domain elements admit continuous representatives. Moreover, the restrictions are closed symmetric operators. It can be proven (see the Appendix) that they both have defect indices $(1,1)$. So, by the von Neumann-Kreĭn theory, they admit a one parameter family of self-adjoint extensions. Their definition and properties are given in the two following propositions, the proof of which is an adaptation of the analogous proof valid for the case of $\mathbb{R}^{3}$ or for the case of bounded domains, and it is reported in detail in the Appendix.
@@
[…]
@@
$(-\Delta_{\alpha,y}^{\textup{D}}-z^{2})^{-1}$ $=(-\Delta^{\textup{D}}-z^{2})^{-1}+(\Gamma_{\alpha,y}^{\textup{D}})^{-1}\langle\overline{G_{z,y}^{\textup{D}}},\cdot\rangle_{L^{2}\left(\mathbb{R}_{+}^{3}\right)}G_{z,y}^{\textup{D}}$ (2.1)
@@
In particular, $\alpha=\infty$ corresponds to $\Delta^{\textup{D}}$.
@@
[…]
@@
$(-\Delta_{\alpha,y}^{\textup{N}}-z^{2})^{-1}$ $=(-\Delta^{\textup{N}}-z^{2})^{-1}+(\Gamma_{\alpha,y}^{\textup{N}})^{-1}\langle\overline{G_{z,y}^{\textup{N}}},\cdot\rangle_{L^{2}\left(\mathbb{R}_{+}^{3}\right)}G_{z,y}^{\textup{N}}$ (2.2)
@@
In particular, $\alpha=\infty$ corresponds to $\Delta^{\textup{N}}$.
@@
The generic domain element of a point interaction is decomposed in the sum of a regular part ($\varphi^{z}$) belonging to the domain of the Laplacian with relevant boundary conditions and a singular part ($qG_{z,y}^{\textup{D}}$ or $qG_{z,y}^{\textup{N}}$) proportional to the Green’s function for the given domain and boundary condition. The decomposition is complemented by a relation linking the regular part at $y$ and the coefficient $q$ of the singular part, playing the role of a further boundary condition at the singular point $y$. For Dirichlet or Neumann case, the boundary conditions at the singularity are given by $\varphi^{z}(y)=\Gamma_{\alpha,y}^{\textup{D}}q$ or $\varphi^{z}(y)=\Gamma_{\alpha,y}^{\textup{N}}q$ respectively. The latter contains the parameter $\alpha$ which specifies the particular self-adjoint extension of the free Laplacian among the whole family of its self-adjoint extensions. This pattern is typical of point interactions in their various incarnations. We add that the physical interpretation of $\alpha$ is related to the scattering length of the interaction.
3. Resonances of point interactions in the half-space
A first result of this work concerns the localization of the continuation resonances for the operators $\Delta_{\alpha,y}^{\textup{D}}$ and $\Delta_{\alpha,y}^{\textup{N}}$. To introduce the concept of resonances for these perturbations of the Laplacian we first recall the corresponding definition for the free Laplacian on the half space. The resolvent $\left(-Delta^{\textup{D}}-z^{2}\right)$ is defined as follows (we consider the Dirichlet case)
$(-\Delta^{\textup{D}}-z^{2})^{-1}:L^{2}(\mathbb{R}_{+}^{3})\to H^{2}(\mathbb{R}_{+}^{3}),\quad\operatorname{Im}{z}>0$
$\left[(-\Delta^{\textup{D}}-z^{2})^{-1}f\right](x)=\int_{\mathbb{R}_{+}^{3}}G_{z,y}^{\textup{D}}(x)f(y)\mathop{}\!\mathrm{d}y.$
This is defined only for $\operatorname{Im}{z}>0$ because the kernel $G_{z,y}^{\textup{D}}(x)$ fails to be in $L^{2}(\mathbb{R}_{+}^{3})$ for $\operatorname{Im}{z}\leq 0$ since it becomes non square integrable in a neighborhood of infinity. If we restrict its domain only to functions with compact support, this behavior of the kernel ceases to be problematic:
$(-\Delta^{\textup{D}}-z^{2})^{-1}:L_{\textup{comp}}^{2}(\mathbb{R}_{+}^{3})\to H_{\textup{loc}}^{2}(\mathbb{R}_{+}^{3}).$ (3.1)
The issue is that $L_{\textup{comp}}^{2}(\mathbb{R}_{+}^{3})$ and $H_{\textup{loc}}^{2}(\mathbb{R}_{+}^{3})$ (or also $L_{\textup{loc}}^{2}(\mathbb{R}_{+}^{3})\supset H_{\textup{loc}}^{2}(\mathbb{R}_{+}^{3})$) are not Banach spaces and to introduce in the simplest way the procedure of holomorphic continuation, a common and useful trick consists in bracketing the resolvent by a smooth cut-off function $\rho$, so obtaining the so called truncated resolvent
$\left[\rho(-\Delta^{\textup{D}}-z^{2})^{-1}\rho f\right](x)=\rho(x)\int_{\Omega}G_{z,y}^{\textup{D}}(x)\rho(y)f(y)\mathop{}\!\mathrm{d}y.$ (3.2)
This is a map between Banach spaces: $\rho(-\Delta^{\textup{D}}-z^{2})^{-1}\rho:L^{2}(\mathbb{R}_{+}^{3})\to L^{2}(\mathbb{R}_{+}^{3})$ for every cut-off function $\rho$ that turns out to be well defined and holomorphic $\forall z\in\mathbb{C}$. So, by definition, the free resolvent can be holomorphically continued to $\mathbb{C}$ if the truncated resolvent can be holomorphically continued to $\mathbb{C}$ as a map $\rho(-\Delta^{\textup{D}}-z^{2})^{-1}\rho:L^{2}(\mathbb{R}_{+}^{3})\to L^{2}(\mathbb{R}_{+}^{3})$.
Now we consider the single point interactions on $\mathbb{R}_{+}^{3}$. We want to carry out the same extension as done for the free Laplacian. Looking at the action of the resolvent in formula Proposition˜2.1 we observe that:
    The first term is simply the free resolvent, which can be holomorphically extended to the whole $\mathbb{C}$ as recalled above.
    In the second term, the projection onto the function $G_{z,y}^{\textup{D}}$, viewed as an operator $L_{\textup{comp}}^{2}(\mathbb{R}_{+}^{3})\to L_{\textup{loc}}^{2}(\mathbb{R}_{+}^{3})$, can be holomorphically extended to all $\mathbb{C}$ in the sense defined above.
    In the second term, the factor $(\Gamma_{\alpha,y}^{\textup{D}}(z))^{-1}$ in front of the projector, is well defined out of the zeroes of $\Gamma_{\alpha,y}^{\textup{D}}(z)$.
The same procedure can be performed for $\Delta_{\alpha,y}^{\textup{N}}$. Summarizing we can state the following proposition and definitions (here $H$ denotes either $\Delta_{\alpha,y}^{\textup{D}}$ or $\Delta_{\alpha,y}^{\textup{N}}$ and $\Gamma_{\alpha,y}$ either $\Gamma_{\alpha,y}^{\textup{D}}$ or $\Gamma_{\alpha,y}^{\textup{N}}$).
Proposition 3.1.
The map
$z\mapsto(H-z^{2})^{-1}:L_{\textup{comp}}^{2}(\mathbb{R}_{+}^{3})\to L_{\textup{loc}}^{2}(\mathbb{R}_{+}^{3})$
extends as a meromorphic family of operators to $z\in\mathbb{C}$.
Definition 3.1.
We call resonances the poles of the meromorphic continuation of $H$, or explicitly, the solutions $z\in\mathbb{C}$ of the equation
$\Gamma_{\alpha,y}(z)=0.$ (3.3)